# FADE: A Dataset for Detecting Falling Objects around Buildings in Video

### v1

@@ V-C Analysis of optical flow based methods in FODB
V-D Visualization Results
Optical flow images. As can be seen in Figure 9, we plot the optical flow images generated by 4 popular optical flow methods (FlowNet 2.0 [102], RAFT [89], PWC-Net [88], and TV-L1 [103]) to visualize the performance of the optical flow in our FADE dataset. The first three groups of optical flow images, estimated by the deep learning based optical flow algorithms [86, 89, 88], are very confusing, resulting in the background and the moving object being difficult to be distinguished. For the last group (i.e. last column) of optical flow images, which are captured by the traditional variational optical flow method [103], the movement of the static background is well estimated, but the motion of the moving objects is also poorly computed.
@@ V-D Visualization Results
Visualization results of representative methods. Visualization results of the representative moving object detection method (MOG2 [42]), image object detection method (Faster R-CNN [78]+FPN [79]), video object detection method (MEGA [100]), and our proposed FADE-Net are shown in Figure 10. We can find that many detection results of MOG2 only include part of the falling object, because MOG2 can only update the model online in the inference process and cannot make use of the rich appearance features of the falling object in the offline training process. Besides, the post-processing of MOG2 includes the opening operation in morphology, which would eliminate the small detection. Thus, some small falling objects on the video frame cannot be detected. The accuracy of Faster R-CNN + FPN is higher than the other two algorithms, and it is also difficult to detect some small falling objects. In addition, the motion blur caused by the fast motion also increases the detection difficulty of the image-based detection model. The accuracy of MEGA is the worst. This method does not learn the appearance features and motion information of the falling object well during the training process, and the lack of non-maximum suppression processing further reduces its accuracy. Our proposed method FADE-Net achieves the best performance in all three video sequences, benefiting from seamlessly integrating generic object detection and motion information in videos. The moving attention module designed by us allows the network to assign adaptive weights to different regions according to the motion information of the image, effectively reducing false detections in the image.
VI Limitation
Generally, our dataset has two main limitations:
The diversity of scenes. Although our dataset includes a large amount of video data and diverse categories, the scenarios can still be further enriched. Consequently, gathering video data in more scenarios to enrich our dataset is a successive future work for FODB.
The baseline method. Our baseline method considers object motion within a short temporal range (less than 1 second). Consequently, exploring to leverage long-term temporal motion information to further boost the detection accuracy is another successive future work for FODB.
VII Conclusion
@@
In this work, we propose a new large-scale video dataset termed FADE for falling object detection around buildings (FODB). It contains 1,881 videos and 164,314 video frames. The videos in FADE include various categories of objects and are captured under diverse scenes, weather conditions, lighting conditions, and video resolutions. Notably, different from the existing MOD datasets, our FADE is the first dataset specialized for FODB. Additionally, we introduce a new baseline method called FADE-Net, which seamlessly integrates motion information capturing and small-sized proposal mining into the detection network. Furthermore, to evaluate the FODB algorithms more effectively, we design an evaluation metric called TRO, which measures the algorithm’s ability to locate the beginning and ending times of falling incidents.

### v3

@@ V-D Analysis of optical flow based methods in FODB
V-E Visualization Results
@@
Visualization results of the representative moving object detection method (MOG2 [45]), image object detection method (Faster R-CNN [80]+FPN [81]), video object detection method (MEGA [101]), and our proposed FADE-Net are shown in Figure 9. Many detection results of MOG2 only capture part of the falling object, as MOG2 applies morphological opening in its post-processing, which tends to eliminate small detections. Thus, some small falling objects on the video frame cannot be detected. The accuracy of Faster R-CNN + FPN is higher than the other two algorithms. However, this method is also difficult to detect some small falling objects. In addition, the motion blur caused by the fast motion also increases the detection difficulty of the image-based detection model. The accuracy of MEGA is the worst. This method does not learn the appearance features and motion information of the falling object well during the training process, and the lack of non-maximum suppression processing further reduces its accuracy. Our proposed method FADE-Net achieves the best performance in all three video sequences, benefiting from seamlessly integrating generic object detection and motion information in videos. In addition, the designed Moving Attention Module enables the network to assign adaptive weights to different regions based on the motion information, effectively reducing false detections in the image.
V-F Effect of incorporating long-term motion information
Although our FADE-Net primarily focuses on utilizing short-term temporal motion information due to the limited computational capabilities of surveillance camera chips, in this section, we also explore the effectiveness of incorporating long-term motion information into our method.
Specifically, we revise our method to incorporate long-term temporal information to experiment with long-term temporal motion data. The revised method adopts a two-stream architecture: one stream captures short-term motion, while the other models long-term motion across the current frame and the preceding five frames. Here, we evaluate the revised model on the FADE dataset using a GTX 1080 GPU to simulate a realistic deployment scenario. The experimental results show that the performance of our original model is comparable to that of the revised model (see Table IV). We attribute this to the high velocity of falling objects, which causes large motion displacements over the long-term window, making it challenging for the model to extract reliable long-term motion cues. Moreover, we can observe from the experimental results that the revised model incurs a certain loss in inference efficiency compared to the original model (FPS 8.9 vs 15.7), which may limit its applicability in real-world falling object detection scenarios. Therefore, we focus on utilizing short-term temporal motion information.
V-G Effect of introducing domain adaptation techniques into our method
Since our falling object detection method is intended for real-world deployment and is expected to operate continuously, it may encounter unseen weather scenarios that were not present during training. To address this, we explore the integration of domain adaptation techniques [108, 109, 110, 111], commonly used in low-level vision tasks [112, 108, 109, 110, 111] such as image deraining [108, 109] and defogging [110, 111] to improve robustness, into our method and evaluate their effectiveness in this context.
Specifically, we design two variants of our method by integrating it with three different domain adaptation strategies, including an adversarial training-based approach using the Gradient Reversal Layer (GRL) [113], as well as a test-time domain adaptation methods, Tent [114]. To better evaluate these variants, we re-partitioned our proposed dataset so that the weather types in the training, validation, and test sets do not overlap. To be specific, the training set contains fair and cloudy conditions, the validation set contains overcast conditions, and the test set contains rainy conditions. Notably, this split is used only for this evaluation.
TABLE IV: Effect of incorporating long-term motion information in the proposed FADE-Net. The better result is indicated in bold.
Method F-measure Precision Recall TRO FPS
Ours w/o long-term 72.03 73.48 70.70 51.75 15.7
Ours w/ long-term 72.08 73.52 70.69 51.77 8.9
The experimental results are presented in Table V above. From the results, we observe that the performance of Ours and Ours w/ GRL is comparable. We believe this is because, although the domain adaptation technique (GRL) encourages the backbone to extract domain-invariant appearance features, in our task where the target objects are extremely small, effectively leveraging motion features is more critical than enhancing appearance features. This observation is further supported by the ablation study, which compares the effectiveness of the motion-focused Moving Attention Module and the appearance-focused Small-Object Mining RPN. In addition, as shown in Table V, our method benefits from the incorporation of the test-time domain adaptation method Tent. However, since Tent operates during inference, it introduces additional computational overhead and reduces the algorithm’s FPS, making it less suitable for real-world deployment. Therefore, considering the overall trade-off between accuracy and inference speed, we adopt “Ours” as the final version.
VI Conclusion
@@
In this work, we propose a new large-scale video dataset termed FADE for falling object detection around buildings (FODB). It contains 2,611 videos and 245,177 video frames. The videos in FADE include various categories of objects and are captured under diverse scenes, weather conditions, lighting conditions, and video resolutions. Notably, different from the existing MOD datasets, our FADE is the first dataset specialized for FODB. Additionally, we introduce a new baseline method called FADE-Net, which seamlessly integrates motion information capturing and small-sized proposal mining into the detection network. Furthermore, to better evaluate the FODB algorithms, we design an evaluation metric called TRO, which measures the algorithm’s ability to locate the beginning and ending times of falling incidents.
@@
TABLE V: Effect of domain adaptation techniques in FADE-Net. We use a GTX 1080 to evaluate the FPS of the algorithms. The better result is indicated in bold.
Method F-measure Precision Recall TRO FPS
Ours 72.08 73.52 70.69 51.77 15.7
Ours w/ GRL [113] 72.07 73.60 70.65 51.78 15.7
Ours w/ Tent [114] 72.15 73.63 70.73 51.80 10.7